OpenStudy (anonymous):

Let theta be an angle in standard position with point ( x, y) on its terminal side. Write the values of the six trigonometric functions of theta for the following values of x, y, and r. x = 0, y = -2, r = 2 sec theta = ?

OpenStudy (anonymous):

secant is the reciprocal of cosine cosine is the first coordinate however, zero has no reciprocal, because you cannot divide by zero

OpenStudy (anonymous):

actually in your case \(\cos(x)=\frac{0}{2}=0\) and \(\sin(x)=\frac{-2}{2}=-1\) but you still cannot divide by 0

OpenStudy (unklerhaukus):